• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name brads

    Log entry time 21:41:07 on February 14, 2007

    Entry number 193210

    This entry is a followup to: 193198

    keyword=BPMs not functioning correctly...

    The BPMs are not functioning correctly at 0.2 or 0.3 uA beam current. At 0.2 uA neither BPM will report values consistently. At 0.3 uA BPM A works but BMP B does not.

    The BPMs worked at these currents prior to today's beam studies, so I assume an IOC got rebooted and/or a gain knob was change that has to be corrected.

    MCC is also telling me that the FFB isn't working at the requested currents and has been disabled (< 0.5 uA so far). Not surprising since BPM-B isn't working.

    I spoke with MCC and they are looking into it some more.